OpenStudy (anonymous):

g(x)=3x h(x)=3x+a find (g -5h) (x) my answer is -12x+20

OpenStudy (anonymous):

so 3x - 5 times h 5(3x+a) = 15x+5a so we have 3x - (15x + 5a) which is -12x -5a

OpenStudy (anonymous):

h(x)= 3x+4

OpenStudy (anonymous):

my mistake on the problem

hero (hero):

g(x) = 3x h(x) = 3x + 4 (g - 5h)(x) = g(x) - 5h(x) = 3x - 5(3x + 4) = 3x - 15x - 5(4) = (3 - 15)x - 20 = -12x - 20

OpenStudy (anonymous):

right thanks man. I had put the problem wrong

hero (hero):

g(x) = 3x h(x) = 3x + 4 (g - 5h)(x) = g(x) - 5h(x) = 3x - 5(3x + 4) = 3x - 15x - 5(4) = (3 - 15)x - 20 = -12x - 20

OpenStudy (anonymous):

its good thanks